the Value M(in dollars)of a laptop computer t years after it was purchased new can be estimated using the function M(t)=800/t+100 where t is greater than or equal to 1. Estimate the laptop's value 4 years after purchase. what does the value fo teh laptop approach as time passes?
Answer by jorel1380(3719)

M=800/t+100
---------------
M=800/4+100=200+100=$300
The value approaches $100 as t grows..